ACC 112 Project 1D Following are independent situations:

Nicholas Ram Corporation have a $2,400,000 "bond issue" dated March 1, 2016 due in 15 years with an annual interest rate of 10%. Interest is payable March 1 and September 1. On August 1, 2016, the bond was sold for $2,513,750 plus accrued interest.

Using the straight-line method, prepare the general journal entries for each of the following:

a)
The issuance of the bond on August 1, 2016.
b)
Payment of the semi-annual interest and the amortization of the premium on September 1, 2016.
c)
Accrual of the interest and the amortization of the premium on December 31, 2016.
d)
Payment of the semi-annual interest and the amortization of the premium on March 1, 2017.


(Credit account titles are automatically indented when the amount is entered. Do not indent manually. Do not use dollar signs ($) when entering amounts. To see comma-formatted numbers reflected in your final answers, you must enter your answers with commas. Round answers to 2 decimal places, e.g. 5,275.25.)


-------------------------------------------------------------------------------------------------------


Stephanie Ram Corporation have a $960,000 "bond issue" dated February 1, 2016 due in 10 years with an annual interest rate of 9%. Interest is payable February 1 and August 1. On April 1, 2016, the bond was sold for $883,300 plus accrued interest.

Using the straight-line method, prepare the general journal entries for each of the following:

a)
The issuance of the bond on April 1, 2016.
b)
Payment of the semi-annual interest and the amortization of the discount on August 1, 2016.
c)
Accrual of the interest and the amortization of the discount on December 31, 2016.
d)
Payment of the semi-annual interest and the amortization of the discount on February 1, 2017.


(Credit account titles are automatically indented when the amount is entered. Do not indent manually. Do not use dollar signs ($) when entering amounts. To see comma-formatted numbers reflected in your final answers, you must enter your answers with commas. Round answers to 2 decimal places, e.g. 5,275.25.)
Answer #1
a) The issuance of the bond on August 1, 2016










DateAccount Titles & ExplanationDebitCredit


Aug. 1Cash$2,578,750




Premium on Bonds payable
$78,750



Bonds Payable
$2,400,000



Bonds Interest payable
$100,000









Interest is accrued for 5 months











b) Payment of the semiannual interest and the amortization of the premium on Sept 1, 2016







Sept. 1Interest expense$20,000




Premium on Bonds payable$2,625




Bonds Interest payable$100,000




Cash
$122,625









Semiannual Interest expense = $2400000 x 10% x 1/2 = $120000


Premium amortization = $78750 / 15 years x 1/2 = $2625 semiannually








c) Accrual of the interest and the amortization of the premium on December 31, 2016







Dec. 31Interest expense$80,000




Premium on Bonds payable$1,750




Interest payable
$81,750









Interest expense for 4 months = $240000/12 x 4 months = $80000


Premium amortization for 4 months = $5250/12 x 4 months = $1750








d) Payment of semiannual interest and the amortization of premium on March 1, 2017







Mar. 1Interest payable$81,750




Interest expense$40,000




Premium on Bonds payable$875




Cash
$122,625









Interest expense for 2 months = $20000 x 2 = $40000



Premium amortization for 2 months = $875










Bonds issued at discount











a) The issuance of the bond on April 1, 2016










DateAccount Titles & ExplanationDebitCredit


April. 1Cash$688,800




Discount on bonds payable$82,600




Bonds Payable
$760,000



Bonds Interest payable
$11,400









Interest accrued for 2 months











b) Payment of the semiannual interest and the amortization of the discount on August 1, 2016







Aug. 1Interest expense$22,800




Bonds Interest payable$11,400




Discount on bonds payable
$4,130



Cash
$30,070









Semiannual Interest expense = $760000 x 9% x 1/2 = $34200


Discount amortization = $82600 / 10 years x 1/2 = $4130 semiannually








c) Accrual of the interest and the amortization of the discount on December 31, 2016







Dec. 31Interest expense$28,500




Discount on bonds payable
$3,442



Interest payable
$25,058









Interest expense for 5 months = $68400/12 x 5 months = $28500


Discount amortization for 5 months = $8260/12 x 5 months = $3442








d) Payment of semiannual interest and the amortization of discount on Feb 1, 2017







Feb. 1Interest payable$25,058




Interest expense$5,700




Discount on bonds payable
688



Cash
$30,070









Interest expense for 1 months = $5700




Discount amortization for 1 months = $688




answered by: Taintal

> This is incorrect

Jahianna Wed, Apr 21, 2021 9:51 PM